# Effects of dark boson mediated feeble interaction between dark matter (DM) and quark matter on $f$-mode oscillation of DM admixed quark stars

### v1

@@ IV Summary and Conclusion
The prominent non-radial $f$-mode frequencies of the DMSQSs are calculated by considering the general relativistic Cowling approximation. By obtaining the $f$-modes as a function of stellar mass, compactness and tidal deformability, we have studied in detail how the presence of DM and its interaction with SQM affect the $f$-mode spectra. We have found that the $f$-mode frequencies of DMSQSs, which are largely populated with massive DM fermions, are higher compared to that of the no-DM scenario (SQSs). We also obtained that the fundamental $f$-mode oscillations are more sensitive to the variations in Fermi momentum of DM ${k_{F}}_{\chi}$, than to the changes in its mass $m_{\chi}$. Thus the presence and interaction of DM with SQM has profound impact on the structural and oscillation properties of the DMSQSs. We conclude that the presence of massive DM of higher density resulted in comparatively less massive, but very compact DMSQSs with larger $f$-mode oscillation frequencies. An empirical relation between the $f$-modes and the average density $\bar{\rho}=\sqrt{\bar{M}/\bar{R}^{3}}$ was also determined and we found that the linear fit for DMSQSs differ significantly from the earlier works. We also estimated relations between the mass-scaled angular frequency $\omega M$ with the compactness $C$. It was observed that $\omega M$ varies universally with $C$ for DMSQSs.
@@ Acknowledgements
L.J.N. acknowledges the Department of Science and Technology, Government of India for the INSPIRE Fellowship. Work of DS is supported by the NRF research Grants (No. 2018R1A5A1025563). Work of AG is supported by the National Research Foundation of Korea, grant funded by Korea Government (MSIT) (RS-2024-00356960).

### v2

@@ IV Summary and Conclusion
The prominent non-radial $f$-mode frequencies of the DMSQSs are calculated by considering the general relativistic Cowling approximation. By obtaining the $f$-modes as a function of stellar mass, compactness and tidal deformability, we have studied in detail how the presence of DM and its interaction with SQM affect the $f$-mode spectra. We have found that the $f$-mode frequencies of DMSQSs, which are largely populated with massive DM fermions, are higher compared to that of the no-DM scenario (SQSs). We also obtained that the fundamental $f$-mode oscillations are more sensitive to the variations in Fermi momentum of DM ${k_{F}}_{\chi}$, than to the changes in its mass $m_{\chi}$. Thus the presence and interaction of DM with SQM has profound impact on the structural and oscillation properties of the DMSQSs. We conclude that the presence of massive DM of higher density resulted in comparatively less massive, but very compact DMSQSs with larger $f$-mode oscillation frequencies. An empirical relation between the $f$-modes and the average density $\bar{\rho}=\sqrt{\bar{M}/\bar{R}^{3}}$ was also determined and we found that the linear fit for DMSQSs differ significantly from the earlier works. We also estimated relations between the mass-scaled angular frequency $\omega M$ with the compactness $C$ and tidal deformability $\Lambda$. It was observed that $\omega M$ varies universally with $C$ for DMSQSs. We also found that for SQS and DMNS, the Cowling approximation has a deviation of about $15-35\%$ when compared with fully relativistic formalism. Further, our investigation reveals that the inclusion of DM reduces the deviation for DMNS, since it makes the star more compact. Similarly, it is expected that DMSQSs may not introduce any further deviation from the results obtained with fully GR conditions. In future, we would like to do a general relativistic treatment of the $f$-mode oscillations in strange quark stars with the presence of DM to understand this point better.
@@ Acknowledgements
We would like to thank the unknown referee, whose suggestions have improved the quality of the manuscript. We would like to thank S. Shirke for providing the DM EoS data used in Fig. 11 for comparison. L.J.N. acknowledges the Department of Science and Technology, Government of India for the INSPIRE Fellowship. Work of DS is supported by the NRF research Grants (No. 2018R1A5A1025563). Work of AG is supported by the National Research Foundation of Korea, grant funded by Korea Government (MSIT) (RS-2024-00356960).